OpenStudy (mendicant_bias):

I'm wondering what approach someone would recommend to solve the following ODE; I don't have to solve it in any particular way.

OpenStudy (mendicant_bias):

\[(1-x^2)y'=1-x^2+y, \ \ \ |x|<1.\]

OpenStudy (daniellelovee):

substitution?

OpenStudy (mendicant_bias):

Alright, cool; I'm reviewing ODE and haven't done this stuff in a while, so I don't remember what that is/how to do it, but I'm going to go read my old text for a moment and then take a shot at it.

OpenStudy (daniellelovee):

ok :) good luck

OpenStudy (mendicant_bias):

Is this the thing involving Homogeneous Equations of a certain order?

OpenStudy (mendicant_bias):

@iambatman , could you help me out on this? I am stuck.

OpenStudy (anonymous):

This ODE is linear

OpenStudy (anonymous):

In particular, \[y'-\frac{1}{1-x^2}y=1\] with integrating factor \[\ln\mu(x)=\int\frac{dx}{x^2-1}=-\frac{1}{2}\ln|x+1|+\frac{1}{2}\ln|x-1|\]

OpenStudy (mendicant_bias):

\[\mu = e^{\int\limits_{}^{}\frac{1}{1-x^2}dx}=e^{1/2 \ln|x+1|+1/2 \ln|x-1|} = \frac{1}{\sqrt{x+1}}+\sqrt{x-1}\]

OpenStudy (mendicant_bias):

Now dealing with that: \[y=\frac{\int\limits_{}^{}\mu(x)f(x)dx}{\mu(x)}=\frac{\int\limits_{}^{}(1/\sqrt{x+1}+\sqrt{x-1})dx}{1/\sqrt{x+1}+\sqrt{x-1}}\]

OpenStudy (mendicant_bias):

\[\int\limits_{}^{}\frac{1}{\sqrt{x+1} \ dx}=2\sqrt{x+1}+C_{1}; \ \int\limits_{}^{}\sqrt{x-1}\ dx=\frac{2(x-1)^{3/2}}{3}+C_{2}.\]

OpenStudy (mendicant_bias):

Answer is this, for reference:

OpenStudy (anonymous):

With some simplifying, \[\mu(x)=\sqrt{\frac{x-1}{x+1}}\] Then distributing gives \[\sqrt{\frac{x-1}{x+1}}y'-\frac{1}{1-x^2}\sqrt{\frac{x-1}{x+1}}y=\sqrt{\frac{x-1}{x+1}}\] Noting that \[\begin{align*}\frac{d}{dx}\sqrt{\frac{x-1}{x+1}}&=\frac{1}{2\sqrt{\frac{x-1}{x+1}}}\frac{x+1-(x-1)}{(x+1)^2}\\\\ &=\frac{(x+1)^{1/2}}{(x-1)^{1/2}(x+1)^2}\\\\ &=\frac{(x-1)^{1/2}}{(x-1)(x+1)^{3/2}}\\\\ &=\frac{(x-1)^{1/2}}{(x^2-1)(x+1)^{1/2}}\\\\ &=-\frac{1}{1-x^2}\sqrt{\frac{x-1}{x+1}} \end{align*}\] then gives \[\frac{d}{dx}\left[\sqrt{\frac{x-1}{x+1}}y\right]=\sqrt{\frac{x-1}{x+1}}\] Integrating both sides, you have \[y=\sqrt{\frac{x+1}{x-1}}{\large \int}\sqrt{\frac{x-1}{x+1}}\,dx\]

OpenStudy (mendicant_bias):

I'm trying to avoid looking at what you've just written at the moment and work out the kinks in my own algebra without spoiling it, but thank you again for the help; I'm retty certain I'll still need to eventually look at what you've written.

OpenStudy (anonymous):

Before you do any looking, there may be a mistake somewhere with my work or with the given solution.

OpenStudy (anonymous):

The given solution is right, so it's definitely something I did...

OpenStudy (mendicant_bias):

Alright, yeah, I had to take a look, but was in the first place confused about the \[\sqrt{\frac{x-1}{x+1}}\]And how exactly you came to that; the way I was working it out, I would end up, moving from the mu value that I had, with an irreducible collection of terms that resembles a factorable/FOILable expression but isn't, sqrt(x^2-1).

OpenStudy (mendicant_bias):

Here's how I'm doing it:

OpenStudy (mendicant_bias):

\[\frac{1}{\sqrt{x+1}}+\sqrt{x-1} \cdot \frac{\sqrt{x+1}}{\sqrt{x+1}} = \frac{1+(x-1)^{1/2}(x+1)^{1/2}}{\sqrt{x+1}}\]

OpenStudy (mendicant_bias):

\[\frac{1+\sqrt{x^2-1}}{\sqrt{x+1}}\]

OpenStudy (anonymous):

From this \[\ln\mu(x)=-\frac{1}{2}\ln|1+x|+\frac{1}{2}\ln|1-x|\] We get the IF \[\begin{align*}\ln\mu(x)&=\frac{1}{2}\ln\left|\frac{1-x}{1+x}\right|\\\\ &=\ln\sqrt{\frac{1-x}{1+x}}\\\\ \mu(x)&=\sqrt{\frac{1-x}{1+x}} \end{align*}\] I had to check quickly with Mathematica; for some reason I get \(\color{red}{(1-x)}\) instead of \((x-1)\) in the numerator (I think this result is preferred anyway, since we have the stipulation \(|x|<1\)).

OpenStudy (mendicant_bias):

Sorry, just laid down for a bit; so above you just used a property of logarithms to combine the arguments, correct?

OpenStudy (anonymous):

This still works though. We have \[\sqrt{\frac{1-x}{1+x}}y'+\frac{1}{x^2-1}\sqrt{\frac{1-x}{1+x}}y=\sqrt{\frac{1-x}{1+x}}\] but \[\begin{align*}\frac{d}{dx}\sqrt{\frac{1-x}{1+x}}&=\frac{1}{2\sqrt{\frac{1-x}{1+x}}}\frac{-(1+x)-(1-x)}{(1+x)^2}\\\\ &=-\frac{(1+x)^{1/2}}{(1-x)^{1/2}(1+x)^2}\\\\ &=\frac{1}{x^2-1}\sqrt{\frac{1-x}{1+x}} \end{align*}\] So, \[\begin{align*}\frac{d}{dx}\left[\sqrt{\frac{1-x}{1+x}}y\right]&=\sqrt{\frac{1-x}{1+x}}\\\\ y&=\sqrt{\frac{1+x}{1-x}}{\large \int}\sqrt{\frac{1-x}{1+x}}\,dx \end{align*}\]

OpenStudy (anonymous):

Yes

OpenStudy (mendicant_bias):

I do not see how differentiating \[\sqrt{\frac{1-x}{1+x}}y\]yields no y terms whatsoever, as in the statement in the second to last line, I'm taking a look to try to figure out why that is. I understand everything else, but not that.

OpenStudy (mendicant_bias):

In fact, I think it's just the phrasing and order of statements, but I understand almost none of what you just wrote. One moment.

OpenStudy (mendicant_bias):

Oh, nevermind, I understand everything else; it's just still that line prior to the last one.

OpenStudy (anonymous):

Maybe some color-coding will help?\[\begin{align*}\frac{d}{dx}\left[\color{red}{\sqrt{\frac{1-x}{1+x}}}\color{blue}{y}\right]&=\color{red}{\frac{d}{dx}\left[\sqrt{\frac{1-x}{1+x}}\right]}\color{blue}y+\underbrace{\color{red}{\sqrt{\frac{1-x}{1+x}}}\color{blue}{\frac{dy}{dx}}}\\\\ &=\underbrace{\color{red}{\sqrt{\frac{1-x}{1+x}}}\color{blue}{y'}}+\color{red}{\frac{1}{x^2-1}\sqrt{\frac{1-x}{1+x}}}\color{blue}y\\\\ &=\text{LHS of ODE}\end{align*}\]

OpenStudy (mendicant_bias):

I'm not trying to be nitpicky, I swear to god, but it's that the first line in your expression, the RHS of that does not equal the RHS of the second to last line; I get the product rule differentiation and all, just supremely confused about it, it's like you just started writing stuff, and that's clearly not what you did, there was intent behind it, lol.

OpenStudy (anonymous):

There are a lot of expressions up there haha, could you be more specific?

OpenStudy (mendicant_bias):

Sure thing, doing that right now, one sec. (1) (2)

OpenStudy (mendicant_bias):

identical LHS, different RHS expressions, primarily the B&W one that makes absolutely no sense to me. At all. Virtually every other expression, gotcha, but differentiating mu*y, I don't understand how differentiating mu*y, you just got a single term, mu. I'd contend that \[\frac{d}{dx}\Bigg[\sqrt{\frac{1-x}{1+x}}y\Bigg] \neq \sqrt{\frac{1-x}{1+x}}\]As stated in the pictured B&W expression.

OpenStudy (mendicant_bias):

\[\frac{d}{dx}(\mu(x)y(x))=\mu(x)f(x)\]

OpenStudy (mendicant_bias):

Is that what's going on?

OpenStudy (mendicant_bias):

I'm guessing it was the last line of math I wrote; in putting the original problem in the proper form, I forgot that the RHS was multiplied by mu, and since f(x) was just 1, it just meant that the RHS was equal to mu.

OpenStudy (anonymous):

Let get back to basics for a moment. A linear ODE has the form \[y'+f(x)y=g(x)\] The integrating factor is \[\mu(x)=\exp\left(\int f(x)\,dx\right)\] so the ODE can be written as \[\begin{align*}\mu(x)y'+\mu(x)f(x)y&=\mu(x)g(x)\\\\ \exp\left(\int f(x)\,dx\right)y'+\exp\left(\int f(x)\,dx\right)f(x)y&=\\\\ \mu(x)\frac{dy}{dx}+\frac{d\mu(x)}{dx}y&=\\\\ \frac{d}{dx}\left[\mu(x)y\right]&= \end{align*}\] Basically what we have here is a rearrangement of the LHS (the application of the IF) so that we can write the LHS as a derivative. From the beginning, we have that the LHS is equivalent to the RHS by virtue of being an ODE. What we've done is to agebraically manipulate the LHS so that we can isolate the factor of \(y\). So, while no \(y\) terms appear on the RHS, we still know that LHS = RHS by the transitive property.

OpenStudy (mendicant_bias):

Yeah, I get how that works...it wasn't the sheer lack of presence of a y term in and of itself, it's that I thought you were just expressing the LHS in a different manner, not equating the LHS and RHS, which, while identical, have pretty different forms/implications are more important there rather than just screwing around and rearranging identities of a single side.

OpenStudy (mendicant_bias):

You were doing \[\frac{d}{dx}\bigg[\mu(x)y(x)\bigg]=\mu(x)f(x),\]not\[\frac{d}{dx}\bigg[\mu(x)y(x)\bigg]=y'(x)\mu(x)+\mu(x)P(x)y(x)\] Despite all three expressions being equivalent.

OpenStudy (mendicant_bias):

Does that make sense?

OpenStudy (anonymous):

That's not what I was doing though. What I did was a logical jump from the modified ODE (with the IF distributed) to the fact that the LHS was a derivative: \[\begin{align*}\sqrt{\frac{1-x}{1+x}}y'+\frac{1}{x^2-1}\sqrt{\frac{1-x}{1+x}}y&=\sqrt{\frac{1-x}{1+x}}\\\\ \frac{d}{dx}\left[\sqrt{\frac{1-x}{1+x}}y\right]&= \end{align*}\] In that post, everything in between was just verification, that \(\dfrac{d}{dx}\sqrt{\dfrac{1-x}{1+x}}\) was indeed the \(\mu(x)f(x)\) term (which it was, and this allows me to go from line 1 to line 2). I hope this addresses your confusion! I have to get to sleep

OpenStudy (mendicant_bias):

I understand the math you did. That's all I have to say at this point. I'm 99% sure we're talking about a virtually identical approach in pedantically different ways. Either way, my blood pressure has skyrocketed. If you think I didn't understand what you did, I don't know what to tell you. Either way, thanks for the help. Goodnight.

OpenStudy (anonymous):

All I did was carry out the general procedure for solving a linear ODE. If you understand the reasoning behind the IF approach, then you should have no problem following along. Like you said, if there's still some cause for confusion, it's likely due to either of us having learned to solve linear equations in a slightly (but fundamentally the same) different method. (If I come across as hostile or condescending, I certainly don't intend to - I usually give askers the benefit of the doubt, but any misunderstanding or confusion puts me on a "let's start from the beginning" track.)

OpenStudy (mendicant_bias):

It's fine, I was considerably more upset than I should have been and nonetheless reacted in the wrong manner. Thank you for helping me.

OpenStudy (anonymous):

No problem. How's that new question coming along?

OpenStudy (mendicant_bias):

I didn't finish it out entirely; due to time constraints, I'm completing most of these problems up to a certain point and keeping a comprehensive record of each problem/how close it is to completion in the form of either embedded pictures of fully worked solutions from my notebook, or links to partial solutions in OpenStudy threads, on some electronic flashcards. I have a system of flashcards whereby I'm not allowed to put a problem with a given solution in a flashcard until I have at least attempted and completed the majority of that problem and have a record of it somewhere; this encourages me to try at least every single one to completion if I'm going to have it ready in that database to prepare for exams. I'll come back to it soon, but for now I'm trying to start on at least every review problem to some meaningful degree, before my next class tomorrow night.